The following is the information from the national income accounts for a hypothetical country:
GDP Rs. 6000.00
Gross Investment 800.00
Net Investment 200.00
Consumption 4000.00
Govt. purchases of goods & services 1100.00
Govt. Budget Surplus 30.00
What is
a) NDP
b) Net exports
c) Govt. taxes minus transfers
c) Disposable personal income
e) Personal Saving.
1
Expert's answer
2017-01-10T09:31:14-0500
GDP: Y = Rs. 6000,
Gross Investment I = 800,
Net Investment NI = 200,
Consumption C = 4000,
Govt. purchases of goods & services G = 1100,
Govt. Budget Surplus BS = T - G = 30.
a) Depreciation = I - NI = 800 - 200 = 600.
NDP = Y - depreciation = 6000 - 600 = 5400.
b) Net exports NX = Y - C - I - G = 6000 - 4000 - 800 - 1100 = 100.
c) Govt. taxes minus transfers T = G + BS = 1100 + 30 = 1130.
c) Disposable personal income Yd = Y - T = 6000 - 1130 = 4870.
e) Personal Saving S = Yd - C = 4870 - 4000 = 870.